# Quantum Vacuum energy as the origin of Gravity

### v2

@@ II.2 An interpretation of the Gibbons-Hawking entropy of de Sitter space
III Induced renormalization group flow for Newton’s constant
III.1 Effective RG flow for Newton’s constant based on the RG flow for $\rho_{\rm vac}$}.
The results of the last section apply whatever $\rho_{\rm vac}$} is. In this section we use the expression (3) to obtain further results. For simplicity of subsequent discourse let us rewrite equations (19),(20) explicitly for $d=3$:
@@ III Induced renormalization group flow for Newton’s constant
In the expression (3) for $\rho_{\rm vac}$}, by dimensional analysis $\mathfrak{g}$ is a dimensionless coupling constant. We assume $m_{\rm z}$ is the physical renormalized mass and ignore its potential RG flow. There are good reasons to expect that $\mathfrak{g}$ has an RG flow, based on the comments in ALrhovac1 ; ALrhovac2 ; LeClairUniverse and equation (6) where $\alpha=e^{2}/4\pi\hbar c$ is the fine structure constant with a known RG flow coming from QED. If we associate $\mathfrak{g}$ with a marginal coupling, $\mathfrak{g}$ could be marginally relevant or irrelevant. Let us assume the following beta function, where $\mu$ is an energy scale, and increasing $\mu$ corresponds to an RG flow to higher energies [^15] The factor of $1/2\pi$ is motivated by standard definitions of fine-structure constants $\alpha$, such as $e^{2}/(4\pi\hbar c)$ for QED. If $\mathfrak{g}=\alpha=\dfrac{g^{2}}{4\pi\hbar c}$ where $g$ is a gauge coupling with beta-function $\beta_{g}=\dfrac{dg}{d\log\mu}=b\dfrac{g^{3}}{16\pi^{2}}$, then $\beta_{\mathfrak{g}}=b\mathfrak{g}^{2}/2\pi$. For pure QED with $N_{f}$ species of Dirac fermions with $g=e$, $\mathfrak{g}_{0}\approx 1/137$ and $b=4N_{f}/3$. For $\lambda\phi^{4}$ theory, if $\mathfrak{g}=\lambda$ then $b=9/8\pi$. See Peskin .
@@
The above beta function is typically just the 1-loop approximation, however we will argue below, based on cosmological data, that the coupling $\mathfrak{g}$ is presently small so that a 1-loop approximation is justified. We assume $\mathfrak{g}>0$ for a positive $\rho_{\rm vac}$} so that we are dealing with de Sitter space. For $\mathfrak{g}>0$, if $b$ is positive the marginal coupling $\mathfrak{g}$ increases at higher energies, namely it is marginally irrelevent, whereas if $b$ is negative $\mathfrak{g}$ flows to zero in the UV, i.e. is asymptotically free. Let $\mathfrak{g}_{0}=\mathfrak{g}(\mu_{0})$ where $\mu_{0}$ is the cosmological energy scale today at time $t_{0}$. Integrating the beta function is straightforward:
@@
The RG parameter $\widehat{b}$ is the primary new constant for applications to cosmology. As we will see, based on cosmological data, we will propose $\widehat{b}>0$, corresponding to a marginally irrelevant coupling, and unless otherwise stated, below $\widehat{b}$ is positive. The pole at $1-\widehat{b}\log(\mu/\mu_{0})=0$ is commonly referred to as a Landau pole, and signifies the theory is not UV complete. It will play an important role below where we will return to addressing its significance.
@@
since we assume $m_{\rm z}$ is already renormalized to the physical mass. Thus $\rho_{\rm vac}$} in (34) should be replaced by ${\rho_{\rm vac}}(\mu)$. In order to express this equation in terms of ${\rho_{\rm vac}}={\rho_{\rm vac}}(\mu_{0})$ today, it is meaningful to incorporate the RG flow into an induced flow for Newton’s constant:
@@
It is meaningful to incorporate the RG flow for $\rho_{\rm vac}$} into an induced flow for $G_{N}$ since we view $G_{N}$ as being determined in the far future where vacuum energy dominates and only afterwards incorporate adding matter and energy as we will do below. If one interprets ${\cal G}(\mu)$ in (39) as an energy scale dependent Newton’s constant, then one sees that for $\widehat{b}>0$, the strength of gravity decreases as one increases the energy scale $\mu$. At extremely low energy scales $\mu$, ${\cal G}(\mu)\to\infty$. The quantity ${\cal H}(t;\mu)$ is interpreted as the solution to Einstein’s equations with a Newton’s constant ${\cal G}(\mu)$ that depends on the energy scale $\mu$. Then based on (34) one has
@@
III.2 Tying the energy scale $\mu$ to the temperature in cosmology
@@
since as we will see below the main features will persist with the addition of matter. As is also standard, we express $a(t)$ in terms of the redshift $z$, $a(t)\equiv 1/{(1+z(t))}$, such that
@@
We thus identify $\mu_{0}=k_{B}T_{0}$, and $z=0$ with $\mu_{0}$. If we identify $G_{N}={\cal G}(\mu_{0})$, then
${\cal G}(\mu)=G_{N}\left(1-\widehat{b}\log(1+z)\right)=G_{N}\left(1+\widehat{b}\log(a)\right).$ (43)
@@
$\left(\dfrac{\dot{a}}{a}\right)=\pm\sqrt{1+\widehat{b}\log a(t)}~\left(\dfrac{8\pi G_{N}\rho_{{\rm vac};0}}{3}\right)^{1/2}$ (44)
where $\rho_{{\rm vac};0}$ is $\rho_{\rm vac}$} at the scale $\mu_{0}$. This additional kind of time dependence does not originate directly from Einstein’s equations but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (42). This leads to a significant and interesting modification of the de Sitter space solution to $a(t)$. Let us first chose the positive sign ++ on the RHS of (44) such that $\dot{a}>0$ in the future. As we will show, choosing the minus sign will be relevant for the far past. The equation (38) can be explicitly integrated, yielding a “inverted gaussian” correction to de Sitter space, which we will simply refer to as gaussian de Sitter space:
@@
Note that as $\widehat{b}\to 0$, $t_{\rm min}\to$-$\infty$ as expected for de Sitter space. In the next section we will approximate $\widehat{b}\approx 0.02$ based on the so-call Hubble tension, which leads to a very large value for $z_{\rm max}$.
Let us now turn to the far past $t\to$-$\infty$. In ordinary de Sitter space, $\lim_{t\to-\infty}a(t)=0$, although it is known this is not a true geometric coordinate singularity. For our gaussian de Sitter space, the solution (45) is formally still valid for $\infty<t<t_{\rm min}$, and thus valid for all times. In fact
@@
To summarize, the solution (45) is valid for all times $\infty<t<\infty$ where $a(t)>a_{\rm min}$. The complete history of such a Universe is that in the far past $t\to$-$\infty$ the scale factor $a(t)$ is infinitely large and starts to compress down to $a_{\rm min}$ which occurs at a time $t_{\rm min}$. Beyond this time, $t>t_{\rm min}$, the Universe expands until the scale factor $a(t)$ is again infinite at $t=+\infty$. This takes an infinite amount of time, and avoids any geometric singularities at $a=0$. There is no “Big Bang” that occurs at time $t_{\rm min}$, rather at this time the Universe smoothly transitions from a compression to an expansion, and $t_{\rm min}$ is just the time when the Universe is hottest. The Big Bang is now better described as a Swing. An picturesque analogy is the harmonic pendulum with the bottom of the swing corresponding to $a_{\rm min}$, and at the present time the Universe is on the up swing. Based on (49), a consistent possibility is that this process then repeats itself, such that Universe is in a sense oscillating about its vacuum like a pendulum harmonic oscillator. At the top of the swing the Universe is at its largest. If this is the case, then the age of this gaussian de Sitter Universe is even more infinite.
IV Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
Thus far we have only considered vacuum energy density $\rho_{\rm vac}$}. We gave a heuristic derivation of Einstein’s equations in this context from the gravitational Casimir effect we formulated above. The Universe also contains radiation and matter, which we view as excitations above the vacuum, which must be included in $T_{\mu\nu}$. Local conservation requires vanishing of the covariant divergence $\nabla^{\mu}T_{\mu\nu}=0$, which is automatic due to the Bianchi identities. Thus, for instance, Newton’s universal law for the gravitational force between two masses is subsumed as a consequence. At this point one must deal with the usual Einstein’s equations except with an energy dependent Newton’s constant ${\cal G}(\mu)$:

### v3

@@ III A reinterpretation of the entropy of de Sitter space and black holes
IV Induced renormalization group flow for Newton’s constant
IV.1 Justifying our formula for $\rho_{\rm vac}$}.
In this subsection we review the analysis that led us to propose the formula (2) for $\rho_{\rm vac}$} in ALrhovac1 ; ALrhovac2 . As motivation, let us simply consider a real scalar field $\phi$ in $d+1$ spacetime dimensions with euclidean action
${\cal S}=\int d^{d+1}x\,\left(\tfrac{1}{2}\partial_{\mu}\phi\partial_{\mu}\phi+V(\phi)\right).$ (32)
Suppose the potential is simply a mass term $V(\phi)=m\phi^{2}/2$ such that spectrum consists of particles of energy $\omega_{\bf{k}}=\sqrt{{\bf{k}}^{2}+m^{2}}$. Treating the free quantized field as a collection of harmonic oscillators with energy $\hbar\omega_{\bf{k}}/2$, naively the vacuum energy density is
$\rho_{\rm vac,cutoff}=\dfrac{s}{2}\int_{|{\bf{k}}|\leq k_{c}}\dfrac{d^{d}{\bf{k}}}{(2\pi)^{d}}\,\sqrt{{\bf{k}}^{2}+m^{2}},$ (33)
where $s=+/$- corresponds to bosons/fermions. Since the above integral is divergent, we have introduced a cut-off $k_{c}$. The leading terms for $k_{c}\gg m$ are
$\rho_{\rm vac,cutoff}=s\begin{cases}\dfrac{k_{c}^{2}}{4\pi}+\dfrac{m^{2}}{4\pi}\,\log(2k_{c}/m)~~~~~~~~~(d=1)\\ \\ \dfrac{k_{c}^{3}}{12\pi}-\dfrac{m^{3}}{12\pi}~~~~~~~~~~~~~~~~~~~~~~~(d=2)\\ \\ \dfrac{k_{c}^{4}}{16\pi^{2}}-\dfrac{m^{4}}{32\pi^{2}}\,\log(2k_{c}/m)~~~~~(d=3)\end{cases}$ (34)
The above calculation is what led Weinberg to state a Cosmological Constant Problem Weinberg . Namely, if $k_{c}$ is taken to be the Planck scale, then the leading term is off by $120$ orders of magnitude. There are two obvious problems with this calculation. First of all, there is no justification for $k_{c}$ to be the Planck scale since we are dealing with a free quantum field theory in Minkowski space with no gravity. Secondly, one is accustomed to ultra-violet divergences in QFT and how to deal with them in order to compute physical quantities. In such a renormalization procedure the leading $k_{c}^{d+1}$ is discarded. However note that for $d+1$ even there is an unavoidable $\log k_{c}$ divergence, unlike $d+1$ odd, and further renormalization is required in order to obtain something finite and independent of the cut-off. However if there are no additional interactions in $V(\phi)$, there is no clear and physical way to remove the $\log k_{c}$ divergence. Based on the above equations (34), the issue of how to deal with the remaining $\log k_{c}$ divergence is essentially identical in $d=1$ and $d=3$ spatial dimensions, thus one can gain insights on the problem from interacting QFT’s in $d=1$ that are exactly solvable, i.e. integrable.
Firstly, one can argue that in $d+1$ spacetime dimensions ${\rho_{\rm vac}}\propto m^{d+1}/g$ where $m$ is a physical mass, and $g$ is an interaction coupling, such that $\rho_{\rm vac}$} diverges as $g\to 0$ reflecting the unavoidable divergence in the free theory displayed in (34). If $m$ is a physical mass, then by dimensional analysis, $g$ is a dimensionless coupling for an interacting theory. That ${\rho_{\rm vac}}\propto 1/g$ implies that it is intrinsically non-perturbative. It turns out this is exactly what occurs for integrable QFT in 2 spacetime dimensions ZamoZamo0 ; ZamoTBA . Namely for all these integrable theories,
${\rho_{\rm vac}}=\dfrac{m_{1}^{2}}{2\mathfrak{g}}~~~~~~~~~~~~~~~~~~~~~~~(d=1)$ (35)
where $m_{1}$ is the mass of the lightest massive particle, and $\mathfrak{g}$ represents a dimensionless coupling, which is a non-perturbative function of the couplings $g$ in $V(\phi)$. The formula (35) can be derived from the thermodynamic Bethe ansatz, which was first formulated by Yang and Yang YangYang , and generalized to relativistic theories by Al. Zamolodchikov ZamoTBA . That $m_{1}$ is the mass of the lightest massive particle follows from the fact that any particle can probe the vacuum to determine $\rho_{\rm vac}$} and the result should be the same, combined with the S-matrix bootstrap which implies that in principle the S-matrix for higher mass particles can be obtained by bootstrapping the lightest mass particle. For integrable theories all of this is well-understood and previously worked out explicitly for many examples. See for instance KlassenMelzer , and Corrigan for the affine Toda theories based on an arbitrary Lie group.
Let us illustrate with the so-called sinh-Gordon model, which is the affine Toda theory based on the Lie group SU(2), with potential
$V(\phi)=2\mu\cosh(\sqrt{8\pi}\,b\,\phi),$ (36)
where $\mu$ sets the energy scale for the mass $m$ of the single particle in the spectrum, and $b$ is a dimensionless coupling. [^15] Not to be confused with $b$ in (3). Here $\mathfrak{g}=-4\pi\sin(\pi\gamma),~~\gamma\equiv b^{2}/(1+b^{2})$, and one finds
${\rho_{\rm vac}}=-\dfrac{m^{2}}{8\pi b^{2}}\left(1+b^{2}+\dfrac{\pi^{2}}{6}b^{4}+\ldots\right).$ (37)
The above result can also be derived in Feynman diagram perturbation theory DestrideVega . The leading term is non-perturbative in the coupling $b$ and the terms in parentheses correspond to perturbative corrections. Indeed one recognizes $\zeta(2)=\pi^{2}/6$ which is commonplace in perturbation theory, and one can easily check the higher orders in $b$ involve $\zeta(2n)$ for integer $n$.
As a warm-up exercise for 4 spacetime dimensions, it is instructive to understand the leading $1/b^{2}$ term in (37) without relying on integrability. To this end we expand the $\cosh$ potential in (35) in powers of $\phi$:
$V(\phi)=2\mu+\dfrac{m^{2}}{2}\phi^{2}+\dfrac{\lambda}{4!}\phi^{4}+{\cal O}(\phi^{6}),~~~~{\rm where}~~~m^{2}=16\pi\mu b^{2},~~~\lambda=128\mu\pi^{2}b^{4}\,.$ (38)
Minimizing the potential $\partial_{\phi}V(\phi)=0$, in addition to a minimum at $\phi=0$ there is a minimum for non-zero $\phi=\overline{\phi}$ where $\overline{\phi}^{2}=-6m^{2}/\lambda$. Ignoring for the moment that this $\overline{\phi}$ is only real for negative $\lambda$, one finds
${\rho_{\rm vac}}=V(\overline{\phi})=2\mu-\dfrac{3}{2}\dfrac{m^{4}}{\lambda}=-\mu=-\dfrac{m^{2}}{16\pi b^{2}}\,$, (39)
where we have used the expressions for $m^{2}$ and $\lambda$ in (38). This agrees with (37) up to a factor of $2$, the discrepancy arising from the fact that we truncated the $\cosh$ potential to order $\phi^{4}$. One can obtain the exact leading term with a similar semi-classical calculation starting from the $\cosh$ potential (36), where the minimum is at $\overline{\phi}=0$ such that $V(\overline{\phi})=2\mu$. One delicate point is that in the thermodynamic Bethe ansatz the scalar particle corresponding to the field $\phi$ must be treated as a fermion since in the high energy limit the S-matrix S=1S=-1. Thus ${\rho_{\rm vac}}=-2\mu$ which equals the leading term in (37). It is important to note that the value of $\rho_{\rm vac}$} in (39) does not require spontaneous symmetry breaking (SSB).
Let us now turn to 4 spacetime dimensions, where one does not have the powerful tools of integrability available. In recent work ALrhovac1 ; ALrhovac2 we formulated a proper computation of $\rho_{\rm vac}$} which renders it finite and well-defined in 4 dimensional Minkowski space and its theoretical basis is the same as in two spacetime dimensions. In ALrhovac1 we showed how to calculate $\rho_{\rm vac}$} from QFT at finite temperature $T$, which parallels the its computation from the thermodynamic Bethe ansatz in $d=1$. Riemann’s zeta function $\zeta(s)$ and its fundamental properties such as the functional equation that relates $\zeta(s)$ to $\zeta(1-s)$ played an essential role in the analysis in ALrhovac1 , and is analogous to modularity in $d=1$. Namely, let $m_{\rm z}$ be the fundamental renormalized scale of physical particle masses where $m_{\rm z}$ represents the lightest mass particle, and $c(r)$ where $r=m_{\rm z}c^{2}/k_{B}T$ be the scaling function such that the free energy density is given by
for some constant coefficients $c_{\rm uv},c_{4}$. [^16] In 2 spacetime dimensions $c_{\rm uv}$ is the effective Virasoro central charge. In 4 spacetime dimensions, $c_{\rm uv}$ is known from conformal field theory if the theory is asymptotically free, as in QCD. Then
${\rho_{\rm vac}}=-c_{4}\pi^{2}m_{\rm z}^{4}/90.$ (41)
This was used in ALrhovac1 to estimate $\rho_{\rm vac}$} for QCD based on finite temperature lattice results where $c_{\rm uv}$ is known since QCD is asymptotically a free conformal field theory.
One can also in principle compute $\rho_{\rm vac}$} directly from the S-matrix at zero temperature using the form-factor bootstrap ALrhovac2 . One advantage of this formulation that all masses are the physical (renormalized) ones once the 2-particle form factor is properly normalized. The form-factor bootstrap relates the two particle form-factor to the zero particle one, which is a 1-point vacuum expectation value of the field in question. As in two spacetime dimensions, we take $m_{\rm z}$ to be the physical mass of the lightest particle, since in principle the S-matrix and form factors for other particles can be obtained by bootstrapping this lightest mass particle. This is where we need Assumption 1 stated in the Introduction, namely that all particle excitations are coupled. In connection with this, in ALrhovac2 we stated the principle of particular democracy, wherein any particle can be used to probe the Vacuum and its energy density, and they should give the same value for $\rho_{\rm vac}$}, and this is known to be true for the $d=1$ models considered there. Applying this to $\langle 0|T_{\mu\nu}|0\rangle$, this leads to a determination of the coupling $\mathfrak{g}$. As usual, defining $\cal T$} from the S-matrix $S=1+i{\cal T}$, then
where $s=(p_{1}+p_{2})^{2}$ is the Mandelstam variable for 2-particle scattering. It was then shown in ALrhovac2 that this leads to (2) which we repeat here
${\rho_{\rm vac}}=\dfrac{3}{4}\dfrac{c^{5}}{\hbar^{3}}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}}\,~~~~~~~~~~(d=3).$ (43)
For general spatial dimension $d$, the factor of $3/4$ equals $d/(d+1)$ which agrees with (35). The derivation of (43) presented in ALrhovac2 from the form-factor bootstrap was new, even in two spacetime dimensions. In this derivation of the above formula for $\rho_{\rm vac}$}, since it is based on the properly normalized 2-particle form factor, $m_{\rm z}$ is the physical renormalized mass.
For integrable theories in $d=1$, the two above definitions of $\rho_{\rm vac}$}, equations (41) and (43), were shown to agree exactly for a wide variety of integrable quantum field theories ALrhovac1 ; ALrhovac2 . Various consistency checks on this formula were made, including that it must vanish for supersymmetric theories and also theories with a fractional supersymmetries where the hamiltonian is in the center of the universal enveloping algebra of the conserved charges. (See ALrhovac1 and references therein.) The coupling $\mathfrak{g}$ can be positive or negative, and $\rho_{\rm vac}$} can even oscillate in sign for a given model as a function of $\mathfrak{g}$, for instance for the sine-Gordon model where it oscillates between positive and negative, and is zero at the (fractional) supersymmetric points. For the sine-Gordon model, there is a special value of the coupling $\mathfrak{g}$ where the fractional supersymmetry is just ${\cal N}=2$ supersymmetry and $\rho_{\rm vac}$} vanishes there as it should ALrhovac1 . In 4 spacetime dimensions we did not rigorously prove that equations (41) and (43) are equivalent definitions of $\rho_{\rm vac}$}, thus (2) should still be viewed as well-motivated but still conjectural. It’s important to note that the above arguments for the formula (2) do not rely on any kind of spontaneous symmetry breaking. For the remainder of the article we simply assume the formula (2).
In 4 spacetime dimensions, we expect that the “coupling” $\mathfrak{g}$ has a renormalization group flow, based in part on the log’s in (34), and we will assume this in the next section. One can motivate this here by considering for instance $\lambda\phi^{4}$ theory where $\mathfrak{g}=\lambda$ to lowest order in perturbation theory. Based on (39), one expects that to leading order ${\rho_{\rm vac}}\propto m_{\rm z}^{4}/\lambda$. It is known that the marginal coupling $\lambda$ has a non-trivial RG flow Peskin . To 1-loop order,
$\mu\partial_{\mu}\lambda=\dfrac{b}{2\pi}\,\lambda^{2},~~~~~~b=\dfrac{9}{8\pi},$ (44)
where increasing the energy scale $\mu$ corresponds to a flow to higher energies. In this model, $\lambda$ grows in the flow to high energies, which is to say it is a marginally irrelevant coupling.
Independent support for the formula (2) comes from Swampland ideas Festina1 ; Festina2 , however this is rather indirect since the arguments are very different as they rely on black holes. By studying a charged particle of mass $m$ in the presence of a black hole, it was argued that
${\rho_{\rm vac}}\leq\dfrac{m^{4}}{2e^{2}}$ (45)
where here $\mathfrak{g}$ is identified with the quantum electrodynamic (QED) fine structure constant $\alpha=e^{2}/(4\pi\hbar c)$. Actual QED will not play a role in this article since we assume the the particle with mass $m_{\rm z}$ is electrically neutral, however, as we will see, it will be instructive to compare couplings and renormalization group parameters with QED, just to check if they are reasonable. Our result (2) is stronger than (45) since it is not an inequality, but it is still consistent with (45).
IV.2 Effective RG flow for Newton’s constant based on the RG flow for $\rho_{\rm vac}$}.
The formulation of a gravitational Casimir effect in Section II is independent of the detailed properties of $\rho_{\rm vac}$}. In this section we use the expression (2) to further develop our dark-universe  model consisting of only vacuum energy. For simplicity of subsequent discourse let us rewrite equations (17),(18) explicitly for $d=3$:
@@ IV Induced renormalization group flow for Newton’s constant
In the expression (2) for $\rho_{\rm vac}$}, by dimensional analysis $\mathfrak{g}$ is a dimensionless coupling constant. Based on the above arguments, $m_{\rm z}$ is the physical renormalized mass of the lightest particle and we thus ignore its potential RG flow. Also based on arguments of the last sub-section, we expect that $\mathfrak{g}$ has an RG flow, based on the comments in ALrhovac1 ; ALrhovac2 ; LeClairUniverse and equation (45) where $\alpha=e^{2}/4\pi\hbar c$ is the fine structure constant with a known RG flow coming from QED. If we associate $\mathfrak{g}$ with a marginal coupling, $\mathfrak{g}$ could be marginally relevant or irrelevant. Let us assume the following beta function, where $\mu$ is an energy scale, and increasing $\mu$ corresponds to an RG flow to higher energies [^17] The factor of $1/2\pi$ is motivated by standard definitions of fine-structure constants $\alpha$, such as $e^{2}/(4\pi\hbar c)$ for QED. If $\mathfrak{g}=\alpha=\dfrac{g^{2}}{4\pi\hbar c}$ where $g$ is a gauge coupling with beta-function $\beta_{g}=\dfrac{dg}{d\log\mu}=b\dfrac{g^{3}}{16\pi^{2}}$, then $\beta_{\mathfrak{g}}=b\mathfrak{g}^{2}/2\pi$. For pure QED with $N_{f}$ species of Dirac fermions, $\mathfrak{g}_{0}\approx 1/137$ and $b=4N_{f}/3$. For $\lambda\phi^{4}$ theory, if $\mathfrak{g}=\lambda$ then $b=9/8\pi$. See Peskin .
@@
The above beta function is typically just the 1-loop approximation, however we will argue below based on cosmological data, that the coupling $\mathfrak{g}$ is presently small so that a 1-loop approximation is justified. Furthermore, in well-understood QFT’s such as QED, the Landau pole discussed below is not removed by higher order corrections to the beta function. We assume $\mathfrak{g}>0$ for a positive $\rho_{\rm vac}$} so that we are dealing with de Sitter space. For $\mathfrak{g}>0$, if $b$ is positive the marginal coupling $\mathfrak{g}$ increases at higher energies, namely it is marginally irrelevent, whereas if $b$ is negative $\mathfrak{g}$ flows to zero in the UV, i.e. is asymptotically free. Let $\mathfrak{g}_{0}=\mathfrak{g}(\mu_{0})$ where $\mu_{0}$ is the cosmological energy scale today at time $t_{0}$. Integrating the beta function is straightforward:
@@
The RG parameter $\widehat{b}$ is the primary new constant for applications to cosmology in our framework. As we will see, based on cosmological data, we will propose $\widehat{b}>0$, corresponding to a marginally irrelevant coupling, and unless otherwise stated, below $\widehat{b}$ is positive.
The pole at $1-\widehat{b}\log(\mu/\mu_{0})=0$ is commonly referred to as a Landau pole, and signifies the theory is not UV complete. It will play an important role below where we will return to addressing its significance. Landau poles are commonplace in elementary particle physics and condensed matter physics, since they are generic to marginally irrelevant perturbations. See for instance LandauPole1 ; LandauPole2 ; LandauPole3 and references therein. Although QED itself is not relevant to this article, let us mention that it also is not UV complete due to a Landau pole, however in this context it occurs at much higher energies than our proposed $z_{\rm max}$, and is thus commonly viewed as a purely academic issue, and is furthermore complicated by chiral symmetry breaking. (See footnote 17).
@@
since we assume $m_{\rm z}$ is already renormalized to the physical mass. Thus $\rho_{\rm vac}$} in (46) should be replaced by ${\rho_{\rm vac}}(\mu)$. In order to express this equation in terms of ${\rho_{\rm vac}}={\rho_{\rm vac}}(\mu_{0})$ today, it is convenient, and meaningful, to incorporate the RG flow into an induced flow for Newton’s constant:
@@
It is meaningful to incorporate the RG flow for $\rho_{\rm vac}$} into an induced flow for $G_{N}$ since we view $G_{N}$ as being determined in the far future where vacuum energy dominates and only afterwards incorporate adding matter and radiation as we will do below. If one interprets ${\cal G}(\mu)$ in (51) as an energy scale dependent Newton’s constant, then one sees that for $\widehat{b}>0$, the strength of gravity decreases as one increases the energy scale $\mu$. At extremely low energy scales $\mu$, ${\cal G}(\mu)\to\infty$. The quantity ${\cal H}(t;\mu)$ is interpreted as the solution to Einstein’s equations with a Newton’s constant ${\cal G}(\mu)$ that depends on the energy scale $\mu$. Then based on (46) one has
@@
V Exact solution to the scale factor $a(t)$ for the solely dark-universe
@@
[…]
@@
This scaling is expected to hold at least post Big Bang Nucleosynthesis, which is roughly a few minutes after the hot big bang. Since our model thus far only consists of vacuum energy density $\rho_{\rm vac}$}, we will assume the equation (54) is valid for all times. We thus identify $\mu_{0}=k_{B}T_{0}$, and $z=0$ with $\mu_{0}$. If we identify $G_{N}={\cal G}(\mu_{0})$, then
${\cal G}(\mu)={\cal G}(z)=G_{N}\left(1-\widehat{b}\log(1+z)\right)=G_{N}\left(1+\widehat{b}\log(a)\right).$ (55)
@@
$\left(\dfrac{\dot{a}}{a}\right)=\pm\sqrt{1+\widehat{b}\log a(t)}~\left(\dfrac{8\pi G_{N}{\rho_{\rm vac}}(\mu_{0})}{3}\right)^{1/2}.$ (56)
This additional kind of time dependence does not originate directly from Einstein’s equations but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (54). By analogy, the classical Maxwell’s equations by themselves cannot account for the well-known energy scale dependence of fine structure constant $\alpha=e^{2}/(4\pi\hbar c)$, which implies it can depend on temperature, and if the temperature is time dependent, such a time dependence is not captured by the time dependence of Maxwell’s equations. This leads to a significant and interesting modification of the de Sitter space solution to $a(t)$. Let us first chose the positive sign ++ on the RHS of (56) such that $\dot{a}>0$ in the future. As we will show, choosing the minus sign will be relevant for the far past, where the solutions match at $a(t)=a_{\rm min}$. The equation (50) can be explicitly integrated, yielding a “inverted gaussian” correction to de Sitter space, which we will simply refer to as gaussian de Sitter space:
@@
Note that as $\widehat{b}\to 0$, $t_{\rm min}\to$-$\infty$ as expected for de Sitter space. In the next section we will approximate $\widehat{b}\approx 0.02$ based on the so-called Hubble tension, which leads to a very large, but finite value for $z_{\rm max}$ (see eq. (66) below.)
Let us now turn to the far past $t\to$-$\infty$. In ordinary de Sitter space, $\lim_{t\to-\infty}a(t)=0$, although it is known this is not a true geometric curvature singularity. For our gaussian de Sitter space, the solution (57) is formally still valid for $\infty<t<t_{\rm min}$, and thus valid for all times. In fact
@@
To summarize, the solution (57) is valid for all times $\infty<t<\infty$ where $a(t)>a_{\rm min}$. The complete history of such a model universe, which we referred to as the dark-universe, is that in the far past $t\to$-$\infty$ the scale factor $a(t)$ is infinitely large and starts to compress down to $a_{\rm min}$ which occurs at a time $t_{\rm min}$. Beyond this time, $t>t_{\rm min}$, the universe expands until the scale factor $a(t)$ is again infinite at $t=+\infty$. This takes an infinite amount of time, and avoids any geometric singularities at $a=0$. There is no “Big Bang” corresponding to scale factor $a(t)=0$. Rather, at time $t_{\rm min}$ the universe smoothly transitions from a compression to an expansion, and $t_{\rm min}$ is just the time when the universe is hottest. The Big Bang is now better described as a Big Swing. Henceforth, by “big bang” we refer to the time $t_{\rm min}$ which is a very hot big bang and not associated with any curvature singularities at $a=0$. A picturesque analogy is the harmonic pendulum with the bottom of the swing corresponding to $a_{\rm min}$, and at the present time the universe is on the up swing. When we add quantum matter and radiation in the next section, these are viewed as excitations above the vacuum, and as the pendulum is on the downswing it accumulates kinetic energy which can excite these states. Based on (61), a consistent possibility is that this process then repeats itself, such that this universe is in a sense oscillating about its vacuum like a pendulum harmonic oscillator about its ground state. At the top of the swing the universe is at its largest. If this is the case, then the age of this gaussian de Sitter universe is even more eternal.
VI Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
Thus far we have only considered vacuum energy density $\rho_{\rm vac}$} which led to the exact solution of the dark-universe presented above. We gave a heuristic derivation of Einstein’s equations in this context from the gravitational Casimir effect we formulated above. The Universe also contains radiation and matter, which we view as excitations above the vacuum, which must be included in $T_{\mu\nu}$. The standard Friedmann equations have three sources, as in (63) below, but without the $(1+\widehat{b}\log a)$ factor. The question naturally arises: should the renormalization properties of $\rho_{\rm vac}$} be kept to the $\Omega_{\Lambda}$ only, i.e. the $\rho_{\rm vac}$} term, or should it be incorporated also in the $\Omega_{\rm m}$ and $\Omega_{\rm rad}$ terms? We take the following point of view, although it perhaps requires more scrutiny, or at least a more complete argument. The dark-universe  is considered a skeleton of our Universe with pure vacuum energy $\rho_{\rm vac}$} and this led to the emergence of $G_{N}$ from $\rho_{\rm vac}$}. Matter and radiation are excitations over the vacuum, however we continue to assume that $G_{N}$ is determined by vacuum energy since we argued that the latter is the origin of Gravity itself. In our analysis of the dark-universe, the renormalization group flow coming from $\rho_{\rm vac}$} was absorbed into an energy scale $\mu$ dependent Newton’s constant $G_{N}$. If the energy scale $\mu$ is fixed, then we add matter and radiation by demanding local energy-momentum conservation, which requires vanishing of the covariant divergence $\nabla^{\mu}T_{\mu\nu}=0$, which is automatic due to the Bianchi identities. Thus, for instance, Newton’s universal law for the gravitational force between two masses at low energies is subsumed as a consequence. At this point one must deal with the usual Einstein’s equations except with an energy dependent Newton’s constant ${\cal G}(\mu)$. Based on this we propose that: